OpenStudy (anonymous):

can we assume no concavities for any function that the power is smaller than x^3? so any function with x^2 would have no concavities?

OpenStudy (amistre64):

is x^2 a line or a curve?

OpenStudy (amistre64):

curves by default have concavity; lines do not

OpenStudy (anonymous):

aren't all x^2 a curve because they are parabolas?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

the reason why I asked the first question is because in order to find concavities we need to find the second derivative right?

OpenStudy (amistre64):

"need" ?? nah .... but if you want to go that route then fine

OpenStudy (anonymous):

haha ok imma keep it simple for myself and still go this route

OpenStudy (anonymous):

but when I go this route... for an example like this: (x^2)-3. the second derivative becomes 2.

OpenStudy (anonymous):

this is where its throwing me off.

OpenStudy (amistre64):

f(x) = x^2 + 6x + 12 f' = 2x + 6 f'' = 2 is a positive value always so this tells us that an +x^2 is always cave up

OpenStudy (amistre64):

the points where the concavity changes would possible at f'' = 0 or undefined

OpenStudy (anonymous):

hmm ok.. so if my second derivative is just a constant, then its showing me that its always that way no matter what the x is right?

OpenStudy (amistre64):

correct; and since the sign demonstrates if its up or down ... +2 is always a cave-up concavity

OpenStudy (anonymous):

so the question is asking me to find intervals of the concavity.

OpenStudy (anonymous):

they gave me my closed intervals of (0<or equal to x <or equal to 2

OpenStudy (anonymous):

so should I just assume that because they gave me those intervals I can say from 2,infinity concave up?

OpenStudy (amistre64):

can you show me a pic of it?

OpenStudy (anonymous):

they never gave me a picture. or are you asking for the picture of the original graph? I can get that from wolfram...

OpenStudy (amistre64):

is the original question asking you for the concavity between 0 and 2 ?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

then its concavity up between 0 and 2; just not too sure how to format that to your solution

OpenStudy (amistre64):

if this had to do with a program; then i was asking if you could take a screenshot of the problem to see if i could understand what sort of format they might want is all

OpenStudy (anonymous):

yeah... it really doesn't have much info haha but thank you! :)